Curvature in Computer Vision

Cheri Shakiban Departament of Mathematics University of St.Thomas St.Paul, MN 55105 USA Email: c9shakiban@stthomas.edu

A new paradigm; the signature curve, is introduced and its application for Object Recognition is discussed.

Euclidean and equi-affine examples are given and a new approach to the numerical approximation of finding the signature curves is presented.

Applications to a variety of fundamental issues in computer vision, including detection of symmetries, visual tracking, and reconstruction of occlusions is discussed and presented by a video tape.